OpenStudy (anonymous):

How do you factor 2x^2-32? Would it be 2(x+4)(x-4)

OpenStudy (whpalmer4):

\[2x^2-32 = 2(x^2-16) = 2(x-4)(x+4)\]You can also check by multiplication:\[2(x+4)(x-4) = (2x+8)(x-4) = 2x^2 - 8x + 8x - 32 = 2x^2-32\checkmark\] Good work!